OpenStudy (anonymous):

sec(x)=-1on interval 0>x>2pi

OpenStudy (anonymous):

\[\sec(x)=1/\cos(x) \implies 1/\cos(x)=-1 \implies \cos(x)=-1 \implies x=\pi\]

OpenStudy (anonymous):

thats the full thing?

OpenStudy (anonymous):

so it is 3.14

OpenStudy (anonymous):

or just pi.

OpenStudy (anonymous):

are the => the greater than or equal to signs or are they showing the next step?

OpenStudy (anonymous):

its an implies. I am replacing sec(x) with 1/cos(x) both of which are equal to -1. Then solving for x.

OpenStudy (anonymous):

so if i were to write this down, showing my work, id leave out the => ?

OpenStudy (anonymous):

If you're asking how to COPY my work, then I'm not answering...

OpenStudy (anonymous):

-.- no, Im just asking if its actually a part of the equation